# New Perspectives on the Polyak Stepsize: Surrogate Functions and Negative Results

### v1

@@ 1 Introduction
In this landscape of stepsize selection strategies, the Polyak stepsize, proposed by Polyak (1969) stands out for its theoretical elegance and convergence properties. Starting from an arbitrary $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$, the Polyak stepsize is defined as
@@ 2 Related Work
In the pionnering work of  Ermol’ev (1966) stepsizes of the form $\eta_{t}\propto\nicefrac{{1}}{{\|\boldsymbol{g}_{t}\|^{2}}}$ were proposed for non-smooth optimization. Despite the many convergence guarantees enabled by  Ermol’ev (1966)’s framework, in  Polyak (1969) it is noted that linear convergence is not possible with such stepsizes. As an alternative, Polyak suggests the stepsize (1), which is shown to converge at favourable rates with non-smooth convex functions, and strongly convex and smooth functions. In fact, contrary to common belief, Polyak (1969) was the first to show linear convergence with a rate comparable to gradient descent in the smooth and strongly convex case. Furthermore, the case where $f^{\star}$ is estimated was also studied, showing convergence to a level set if $f^{\star}$ is overestimated, and convergence to a neighboorhood if it is underestimated. The Polyak stepsize has since been extended and studied across several applications and domains.
Non-smooth convex. In non-smooth convex optimization, several schemes have been developed to estimate $f^{\star}$ on the fly (Brännlund, 1993; Bertsekas, 1999; Goffin and Kiwiel, 1999; Sherali et al., 2000; Nedic and Bertsekas, 2001; Kiwiel, 2004). In the finite-sum case with interpolation, (1) and variants have been studied as an incremental subgradient method (Nedic and Bertsekas, 2001; Kiwiel, 2004).
Non-expansive operators. In the context of non-expansive operators, the update (1) has also been studied as a special case of the subgradient projector (Bauschke and Combettes, 2017; Cegielski, 2012; Combettes, 2009). Where it can be shown that subgradient descent with $\eta_{t}=\nicefrac{{(f(x)-c)_{+}}}{{\|\boldsymbol{g}_{t}\|^{2}}}$ is a quasi-firmly non-expansive operator [^1] An operator $T$ is quasi-firmly non-expansive if for all fixed points $\boldsymbol{x}^{\star}$, $\|T(\boldsymbol{x})-\boldsymbol{x}^{\star}\|^{2}\leq\|\boldsymbol{x}-\boldsymbol{x}^{\star}\|^{2}-\|T(\boldsymbol{x})-\boldsymbol{x}^{\star}\|^{2}$. Note that a quasi-firmly non-expansive operators are also referred to as cutters. if $f$ is convex and $c\geq f^{\star}$, and $\boldsymbol{x}\to\boldsymbol{x}^{\star}$ if $f$ is continuous (Bauschke and Combettes, 2017). Moreover, in the finite-sum setting, interpolation can also be viewed as iterating different quasi-firmly non-expansive operators with a common fixed point. For example, applying a subgradient projector sequentially (i.e., cycling through the different component functions) in a way such that each function eventually gets visited guarantees that $\boldsymbol{x}_{t}\to\{f(\boldsymbol{y})\leq c\}$ (Cegielski, 2012, Example 5.9.7). For a survey on this topic see Censor (1984).
Deterministic. In modern optimization, (1) has been shown to achieve similar rates to gradient descent in various common assumptions (e.g., Lipschitz, smoothness and strong convexity) (Hazan and Kakade, 2019), and more recently with other assumptions such as weakly convex functions (Davis et al., 2018), $(L_{0},L_{1})$-smooth functions (Takezawa et al., 2024; Gorbunov et al., 2025), and directional smoothness (Mishkin et al., 2024).
@@
Neighbourhood of Convergence. For SPSmax, Loizou et al. (2021) proved that the suboptimality gap is only guaranteed to shrink up to a factor that depends on the loss themselves. As we will explain in Section 5.1, this is equivalent to the guarantees in online learning where the regret is proportional to the cumulative loss of the competitor, typically denoted by $L^{\star}$. Hence, these kind of guarantees are usually called in online learning $L^{\star}$ bounds (see, e.g., Orabona, 2019, Section 4.2).
@@ 3 Definitions and Notation
[…]
@@
Note that $f$ has a sharp minimum then it is not differentiable at $\boldsymbol{x}^{\ast}$ (Polyak, 1987) and if the function is also convex and $G$-Lipschitz we immediately have $G\geq s$.
@@
$\|\boldsymbol{g}\|^{2}\leq 2L(f(\boldsymbol{x})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x})),\ \forall\boldsymbol{x}\in\partial f(\boldsymbol{x})~{}.$
It is known that $L$-smooth functions are $L$-self-bounded (see, e.g., Lemma 4 in Li and Orabona, 2019), but this definition is strictly weaker because it does not assume differentiability.
@@ 4 Polyak Stepsize is Gradient Descent on a Surrogate Function
Instead of viewing the Polyak stepsize (1) with respect to $f$ we propose to view it equivalently as a subgradient method with respect to $\phi$. By the chain rule of subgradients (Bauschke and Combettes, 2017)[Corollary 16.72], subgradient descent with the Polyak stepsize (1) is equivalent to subgradient descent on $\phi$ with stepsize $\eta_{t}=\frac{1}{\|\boldsymbol{g}_{t}\|^{2}}$. This perspective may seem superfluous, howerever, we will show that $\frac{1}{\|\boldsymbol{g}_{t}\|^{2}}$ is strongly related to a certain notion of local curvature of $\phi$, local star upper curvature.
@@
We say that a function $f$ with minimizer $\boldsymbol{x}^{\star}$ has $\lambda_{\boldsymbol{y}}$-local star upper curvature (LSUC) around $\boldsymbol{y}$ if for all $\boldsymbol{y}$ there exists $\lambda_{\boldsymbol{y}}>0$ such that
@@
Note that the definition implies the convexity of $f$ since we assume the existence of a subgradient. [^2] If $\boldsymbol{g}_{t}$ is not a subgradient but a directional derivative then $f$ would be guaranteed to be star-convex (Joulani et al., 2020). It is also immediate to show that convex $L$-smooth functions are also $L$-LSUC. Indeed, for convex $L$-smooth functions we have that (Nesterov, 2004, Theorem 2.1.5)
@@
So, it is enough to set $\boldsymbol{x}=\boldsymbol{x}^{\star}$ to obtain the above definition. However, the inclusion is strict, because there exist functions that are not smooth and still satisfy the above definition. For example, one can easily verify that $f(x)=|x+2|+\frac{x^{2}}{2}$ is $LSUC$ but not differentiable, so not smooth.
Finally, if the star-upper-curvature holds globally, i.e., there exists $0<\lambda<\lambda_{\boldsymbol{y}}$ for all $\boldsymbol{y}$, then we can show that this condition is equivalent to the upper quadratic growth condition in Guille-Escuret et al. (2021). This observation was first made by Goujaud et al. (2022, Theorem 2.6), we include the precise statement and proof in the Appendix B.
@@
This theorem tells us that, regardless of the curvature of $f$, we can always construct the function $\phi$ that is locally curved. It is well-known that for $L$-smooth functions one can use the stepsize $\eta=\frac{1}{L}$ and achieve a rate between $\mathcal{O}(1/T)$ and a linear one, depending on the presence of strong convexity. Here, we show a similar result: GD can use stepsizes that depend on the local star upper curvature in all cases. Note however, unlike GD with a constant stepsize and smoothness, we do not have a descent lemma with $\phi$. Indeed this is not possible as it would guarantee a $\mathcal{O}(1/T)$ rate for the last iterate which was shown to be impossible by Goujaud et al. (2022) for $QG+(L)$ functions (i.e., L-LUSC functions).
@@
Let $\phi$ convex and define $\boldsymbol{x}^{\star}\in\mathop{\mathrm{argmin}}_{\boldsymbol{x}}\ \phi(\boldsymbol{x})$. Assume $f$ to be $\lambda_{\boldsymbol{x}}$-LSUC around any point $\boldsymbol{x}$. Then, using GD with stepsizes $\eta_{t}=\frac{1}{\lambda_{\boldsymbol{x}_{t}}}$ guarantees
@@
From the classic one-step analysis of GD (see, e.g., Orabona, 2019), we have
$\eta_{t}\langle\boldsymbol{g}_{\boldsymbol{x}_{t}},\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\rangle=\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}-\frac{1}{2}\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|^{2}+\frac{\eta_{t}^{2}}{2}\|\boldsymbol{g}_{t}\|^{2}~{}.$
Now, we use the fact that $\phi$ is LSUC and the definition of $\eta_{t}$ we obtain the stated bound. Summing from $t=1$ to $T$ and discarding the negative term on the right hand side concludes the proof. ∎
@@
This theorem and Lemma 1 do not give us a rate, however we can immediately observe that if $\sum_{t}\eta_{t}=+\infty$ we also have $\phi(\boldsymbol{x}_{t})\to 0$, that is, the last iterate will converge.
@@
Similarly, it is equally easy to obtain rates for Hölder-smooth functions.
@@
$\frac{s^{2}}{G^{2}}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}\leq s^{2}\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}\leq\eta_{t}(\phi(\boldsymbol{x}_{t})-\phi(\boldsymbol{x}^{\star}))\leq\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}-\frac{1}{2}\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|^{2}~{}.$
@@ 5 Generalizing the Polyak Stepsize: More Surrogates and Stochastic Setting
where the inequality is due to the convexity of $h$ and the fact that $h(\boldsymbol{x})\geq 0$. Setting $\boldsymbol{u}\in\mathop{\mathrm{argmin}}_{\boldsymbol{x}}\ \psi(\boldsymbol{x})$, we have the stated bound. ∎
@@
The last two lemmas tells us that the properties of the surrogate functions breaks if $\psi(\boldsymbol{x}^{\star})\neq 0$. Hence, we will not be able to prove convergence results, but only that, for example, the suboptimality gap will converge up to a floor that depends on $\psi(\boldsymbol{x}^{\star})$. However, in Section 6 we will show that this is not an artifact of the proof. Indeed, we can construct simple one-dimensional functions where the generalized Polyak stepsize does not converge.
@@ 5.1 Stochastic Approximation Setting
0:  $h:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$
@@
Here, starting from ALI-G (Berrada et al., 2020) and SPSmax (Loizou et al., 2021) that use the idea of limiting the stepsizes, we propose a generalized Polyak stepsize algorithm. The proof is in Appendix C.
@@
  • SPSmax (Loizou et al., 2021): $h(\boldsymbol{x},\boldsymbol{\xi})=f(\boldsymbol{x},\boldsymbol{\xi})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x},\boldsymbol{\xi})$, so $H(\boldsymbol{x}^{\star})=\mathbb{E}[f(\boldsymbol{x}^{\star},\boldsymbol{\xi})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x},\boldsymbol{\xi})]$.
@@
  • SPS+ (Garrigos et al., 2023): $h(\boldsymbol{x},\boldsymbol{\xi})=(f(\boldsymbol{x},\boldsymbol{\xi})-f(\boldsymbol{x}^{\star},\boldsymbol{\xi}))_{+}$. In this case, $H(\boldsymbol{x}^{\star})=0$ so we can also safely set $\gamma=\infty$. Moreover, $H(\boldsymbol{x})\geq F(\boldsymbol{x})-F(\boldsymbol{x}^{\star})$, hence any bound on $H(\boldsymbol{x})$ translates to a bound on the suboptimality gap.
@@
If $h(\boldsymbol{x}^{\star},\boldsymbol{\xi}_{t})=0$ for all $t$, then $\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|\leq\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|$. Hence, in this case we only need to consider all the properties of $h$ in the bounded domain $\{\boldsymbol{x}\in\mathbb{R}^{d}:\|\boldsymbol{x}-\boldsymbol{x}^{\star}\|\leq\|\boldsymbol{x}_{1}-\boldsymbol{x}^{\star}\|\}$. This is well-known via the subgradient projector perspective, as $\boldsymbol{x}_{t}$ is guaranteed to approach the set $\bigcap_{\boldsymbol{\xi}}\{\boldsymbol{x}:h(\boldsymbol{x},\xi)=0\}$ at each iteration if it is non-empty. This property was observed in Gower et al. (2025) for SPS+ (Garrigos et al., 2023), but here it holds more generally. For example, $h(\boldsymbol{x},\boldsymbol{\xi})=(f(\boldsymbol{x},\boldsymbol{\xi})-a)_{+}$, where $a\geq\sup_{\xi}\min_{\boldsymbol{x}}f(\boldsymbol{x},\boldsymbol{\xi})$, would also have no neighbourhood of convergence and satisfies the assumption of the theorem.
@@
In the proof of Theorem 3, if one stops before taking expectations, one obtains a regret guarantee on a sequence of arbitrary losses $h(\boldsymbol{x},\boldsymbol{\xi}_{t})$. Such regret scales as the sum of the loss in $\boldsymbol{x}^{\star}$. This is exactly the $L^{\star}$ bound that we mentioned in Section 2. Indeed, this kind of updates and guarantees were already obtained in the online learning literature for the special case of linear predictors by the Passive-Aggressive family of algorithms (Crammer et al., 2006).
Besides covering a number of previous algorithmic variants, we also extend the previous known guarantees. In particular, Loizou et al. (2021) did not prove a guarantee for the non-smooth setting. So, our result proves for the first time that SPSmax  is adaptive to the entire range of upper curvature of the function, from Lipschitz to smooth functions. In Appendix D we also show additional results. Moreover, the second result in the smooth case is new, and it allows to recover the SGD guarantee on $H$ when $\gamma$ is sufficiently small. For example, we include a precise statement for SPS+ when $f(\boldsymbol{x},\boldsymbol{\xi})$ is $L$-self-bounded, that recovers the guarantee in Gower et al. (2025, Corollary 2.3).
6 Neighbourhood of Convergence and Instability of Polyak Stepsize

### v2

@@ 1 Introduction
In this landscape of stepsize selection strategies, the Polyak stepsize, proposed by Polyak [36] stands out for its theoretical elegance and convergence properties. Starting from an arbitrary $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$, the Polyak stepsize is defined as
@@ 2 Related Work
In the pionnering work of  Ermol’ev [14] stepsizes of the form $\eta_{t}\propto\nicefrac{{1}}{{\|\boldsymbol{g}_{t}\|^{2}}}$ were proposed for non-smooth optimization. Despite the many convergence guarantees enabled by  Ermol’ev [14]’s framework, in  Polyak [36] it is noted that linear convergence is not possible with such stepsizes. As an alternative, Polyak suggests the stepsize (1), which is shown to converge at favourable rates with non-smooth convex functions, and strongly convex and smooth functions. In fact, contrary to common belief, Polyak [36] was the first to show linear convergence with a rate comparable to gradient descent in the smooth and strongly convex case. Furthermore, the case where $f^{\star}$ is estimated was also studied, showing convergence to a level set if $f^{\star}$ is overestimated, and best-iterate convergence to a neighboorhood if it is underestimated. The Polyak stepsize has since been extended and studied across several applications and domains.
Non-smooth convex. In non-smooth convex optimization, several schemes have been developed to estimate $f^{\star}$ on the fly [7, 6, 16, 41, 31, 27]. In the finite-sum case with interpolation, (1) and variants have been studied as an incremental subgradient method [31, 27].
Non-expansive operators. In the context of non-expansive operators, the update (1) has also been studied as a special case of the subgradient projector [2, 8, 10]; where it can be shown that subgradient descent with $\eta_{t}=\nicefrac{{(f(x)-c)_{+}}}{{\|\boldsymbol{g}_{t}\|^{2}}}$ is a quasi-firmly non-expansive operator [^1] An operator $T$ is quasi-firmly non-expansive if for all fixed points $\boldsymbol{x}^{\star}$, $\|T(\boldsymbol{x})-\boldsymbol{x}^{\star}\|^{2}\leq\|\boldsymbol{x}-\boldsymbol{x}^{\star}\|^{2}-\|T(\boldsymbol{x})-\boldsymbol{x}\|^{2}$. Note that a quasi-firmly non-expansive operators are also referred to as cutters. if $f$ is convex and $c\geq f^{\star}$, and $\boldsymbol{x}\to\boldsymbol{x}^{\star}$ if $f$ is continuous [2]. Moreover, in the finite-sum setting, interpolation can also be viewed as iterating different quasi-firmly non-expansive operators with a common fixed point. For example, applying a subgradient projector sequentially (i.e., cycling through the different component functions) in a way such that each function eventually gets visited guarantees that $\boldsymbol{x}_{t}\to\{f(\boldsymbol{y})\leq c\}$ [8, Example 5.9.7]. For a survey on this topic see Censor [9].
Deterministic. In modern optimization, (1) has been shown to achieve similar rates to gradient descent in various common assumptions (e.g., Lipschitz, smoothness and strong convexity) [22], and more recently with other assumptions such as weakly convex functions [13], $(L_{0},L_{1})$-smooth functions [42, 17], and directional smoothness [30].
@@
Neighbourhood of Convergence. For SPSmax, Loizou et al. [29] proved that the suboptimality gap is only guaranteed to shrink up to a factor that depends on the loss themselves. As we will explain in Section 5.1, this is equivalent to the guarantees in online learning where the regret is proportional to the cumulative loss of the competitor, typically denoted by $L^{\star}$. Hence, these kind of guarantees are usually called in online learning $L^{\star}$ bounds [see, e.g., 34, Section 4.2].
Surrogates. Gower et al. [19] show that the Polyak stepsize is equivalent to online gradient descent with fixed stepsize on a sequence of adversarially chosen self-bounded surrogate losses. Differently from our framework, the adversarial nature of their losses does not allow to show that the algorithm is minimizing a fixed function. In comparison, our surrogate approach considers a fixed surrogate loss with local smoothness, where the Polyak stepsize is chosen to be the inverse of the local smoothness.
@@ 3 Definitions and Notation
We will use the following notation and definitions. All the norms in this paper are L2 norms and will be denoted by $\|\cdot\|$. For a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define a subgradient of $f$ in $\boldsymbol{x}\in\mathbb{R}^{d}$ as a vector $\boldsymbol{g}\in\mathbb{R}^{d}$ that satisfies $f(\boldsymbol{y})\geq f(\boldsymbol{x})+\langle\boldsymbol{g},\boldsymbol{y}-\boldsymbol{x}\rangle,\ \forall\boldsymbol{y}\in\mathbb{R}^{d}$. We denote the set of subgradients of $f$ in $\boldsymbol{x}$ by $\partial f(\boldsymbol{x})$. For a differentiable function we have that $\partial f(\boldsymbol{x})=\{\nabla f(\boldsymbol{x})\}$. A function $f:V\rightarrow\mathbb{R}$, differentiable in an open set containing $V$, is $L$-smooth w.r.t. $\|\cdot\|$ if $f(\boldsymbol{y})\leq f(\boldsymbol{x})+\langle\nabla f(\boldsymbol{x}),\boldsymbol{y}-\boldsymbol{x}\rangle+\frac{L}{2}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}$ for all $\boldsymbol{x},\boldsymbol{y}\in V$.
@@
Note that if $f$ has a sharp minimum then it is not differentiable at $\boldsymbol{x}^{\ast}$ [37] and if the function is also convex and $G$-Lipschitz we immediately have $G\geq s$.
@@
$\|\boldsymbol{g}\|^{2}\leq 2L(f(\boldsymbol{x})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x})),\ \forall\boldsymbol{g}\in\partial f(\boldsymbol{x})~.$
It is known that $L$-smooth functions are $L$-self-bounded [see, e.g., Lemma 4 in 28], but this definition is strictly weaker because it does not assume differentiability.
@@ 4 Polyak Stepsize is Gradient Descent on a Surrogate Function
Instead of viewing the Polyak stepsize (1) with respect to $f$ we propose to view it equivalently as a subgradient method with respect to $\phi$. By the chain rule of subgradients [2][Corollary 16.72], subgradient descent with the Polyak stepsize (1) is equivalent to subgradient descent on $\phi$ with stepsize $\eta_{t}=\frac{1}{\|\boldsymbol{g}_{t}\|^{2}}$. This perspective may seem superfluous, howerever, we will show that $\frac{1}{\|\boldsymbol{g}_{t}\|^{2}}$ is strongly related to a certain notion of local curvature of $\phi$, local star upper curvature.
Figure 1: The function $f(x)=|x+2|+\frac{x^{2}}{2}$ is non-smooth but is $2$-LSUC as demonstrated by the blue curve, $f(x^{\star})-\langle\boldsymbol{g},x^{\star}-x\rangle-\nicefrac{{1}}{{4}}\|\boldsymbol{g}\|^{2}$, being larger than $f(x)$ for all $x$ and $\boldsymbol{g}\in\partial f(x)$. Similarly, $f$ is self-bounded but with the larger constant $L=9$.
@@
We say that a function $f$ with minimizer $\boldsymbol{x}^{\star}$ has $\lambda_{\boldsymbol{y}}$-local star upper curvature (LSUC) around $\boldsymbol{y}$ if there exists $\lambda_{\boldsymbol{y}}>0$ such that
@@
Note that if the function is LSUC everywhere, then it must be convex since we assume the existence of a subgradient. [^2] If $\boldsymbol{g}_{t}$ is not a subgradient but a directional derivative then $f$ would be guaranteed to be star-convex [26]. It is also immediate to show that convex $L$-smooth functions are also $L$-LSUC. Indeed, for convex $L$-smooth functions we have that [32, Theorem 2.1.5]
@@
So, it is enough to set $\boldsymbol{x}=\boldsymbol{x}^{\star}$ to obtain the above definition. However, the inclusion is strict, because there exist functions that are not smooth and still satisfy the above definition. For example, as shown in Figure 1, one can easily verify that $f(x)=|x+2|+\frac{x^{2}}{2}$ is $2$-LSUC and $9$-self-bounded but not differentiable x=2x=-2, hence it is not smooth.
Finally, if the star-upper-curvature holds globally, i.e., there exists $0<\lambda<\lambda_{\boldsymbol{y}}$ for all $\boldsymbol{y}$, then we can show that this condition is equivalent to the upper quadratic growth condition in Guille-Escuret et al. [21]. [^3] A function $f$ satisfies the $\mu$-quadratic growth condition if $f(\boldsymbol{x})-f(\boldsymbol{x}^{\star})\geq\nicefrac{{\mu}}{{2}}\|\boldsymbol{x}-\boldsymbol{x}^{\star}\|^{2}$. This observation was first made by Goujaud et al. [18, Theorem 2.6], we include the precise statement and proof in the Appendix B.
@@
This theorem tells us that, regardless of the curvature of $f$, we can always construct the function $\phi$ that is locally curved. It is well-known that for $L$-smooth functions one can use the stepsize $\eta=\frac{1}{L}$ and achieve a rate between $\mathcal{O}(1/T)$ and a linear one, depending on the presence of strong convexity. Here, we show a similar result: GD can use stepsizes that depend on the local star upper curvature in all cases. Note however, unlike GD with a constant stepsize and smoothness, we do not have a descent lemma with $\phi$. Indeed this is not possible as it would guarantee a $\mathcal{O}(1/T)$ rate for the last iterate which was shown to be impossible by Goujaud et al. [18] for $QG+(L)$ functions (i.e., L-LUSC functions).
@@
Let $\phi$ convex and define $\boldsymbol{x}^{\star}\in\mathop{\mathrm{argmin}}_{\boldsymbol{x}}\ \phi(\boldsymbol{x})$. Assume $\phi$ to be $\lambda_{\boldsymbol{x}}$-LSUC around any point $\boldsymbol{x}$. Then, using subgradient descent with stepsizes $\eta_{t}=\frac{1}{\lambda_{\boldsymbol{x}_{t}}}$ guarantees
@@
From the classic one-step analysis of GD [see, e.g., 34], we have
$\eta_{t}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\rangle=\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}-\frac{1}{2}\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|^{2}+\frac{\eta_{t}^{2}}{2}\|\boldsymbol{g}_{t}\|^{2},$
for any $\boldsymbol{g}_{\boldsymbol{x}_{t}}\in\partial\phi(\boldsymbol{x}_{t})$. Now, we use the fact that $\phi$ is LSUC and the definition of $\eta_{t}$ we obtain the stated bound. Summing from $t=1$ to $T$ and discarding the negative term on the right hand side concludes the proof. ∎
@@
This theorem and Lemma 1 do not give us a rate, however we can immediately observe that if $\sum_{t}\eta_{t}=+\infty$ we also have $\phi(\boldsymbol{x}_{t})\to 0$, implying convergence of the last iterate (see Remark 6 for more details).
@@
Similarly, it is equally easy to obtain rates for Hölder-smooth functions, see Theorem 7 for more details.
@@
$\frac{s^{2}}{G^{2}}\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}\leq s^{2}\eta_{t}\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}\leq\eta_{t}(\phi(\boldsymbol{x}_{t})-\phi(\boldsymbol{x}^{\star}))\leq\frac{1}{2}\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|^{2}-\frac{1}{2}\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|^{2}~.$
@@ 5 Generalizing the Polyak Stepsize: More Surrogates and Stochastic Setting
where the inequality is due to the convexity of $h$ and the fact that $h(\boldsymbol{x})\geq 0$. Setting $\boldsymbol{u}=\boldsymbol{x}^{\star}$, we have the stated bound. ∎
@@
The last two lemmas tell us that the properties of the surrogate functions breaks if $\psi(\boldsymbol{x}^{\star})\neq 0$. Hence, we will not be able to prove convergence results, but only that, for example, the suboptimality gap will converge up to a floor that depends on $\psi(\boldsymbol{x}^{\star})$. However, in Section 6 we will show that this is not an artifact of the proof. Indeed, we can construct simple one-dimensional functions where the generalized Polyak stepsize does not converge.
@@ 5.1 Stochastic Approximation Setting
[…]
@@
Here, starting from ALI-G [5] and SPSmax [29] that use the idea of limiting the stepsizes, we propose a generalized Polyak stepsize algorithm. The proof is in Appendix C.
@@
Let $h:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}_{\geq 0}$ be convex in its first argument. Denote by $H(\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\xi}\sim D}[h(\boldsymbol{x},\boldsymbol{\xi})]$. Then, setting $\eta_{t}=\min\left(\frac{1}{\|\boldsymbol{g}_{t}\|^{2}},\frac{\gamma}{h(\boldsymbol{x}_{t},\boldsymbol{\xi}_{t})}\right)$ in Algorithm 1, we have
@@
  • SPSmax [29]: $h(\boldsymbol{x},\boldsymbol{\xi})=f(\boldsymbol{x},\boldsymbol{\xi})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x},\boldsymbol{\xi})$, so $H(\boldsymbol{x}^{\star})=\mathbb{E}[f(\boldsymbol{x}^{\star},\boldsymbol{\xi})-\inf_{\boldsymbol{x}}\ f(\boldsymbol{x},\boldsymbol{\xi})]$.
@@
  • SPS+ [15]: $h(\boldsymbol{x},\boldsymbol{\xi})=(f(\boldsymbol{x},\boldsymbol{\xi})-f(\boldsymbol{x}^{\star},\boldsymbol{\xi}))_{+}$. In this case, $H(\boldsymbol{x}^{\star})=0$ so we can also safely set $\gamma=\infty$. Moreover, $H(\boldsymbol{x})\geq F(\boldsymbol{x})-F(\boldsymbol{x}^{\star})$, hence any bound on $H(\boldsymbol{x})$ translates to a bound on the suboptimality gap.
@@
If $h(\boldsymbol{x}^{\star},\boldsymbol{\xi}_{t})=0$ for all $t$, then $\|\boldsymbol{x}_{t+1}-\boldsymbol{x}^{\star}\|\leq\|\boldsymbol{x}_{t}-\boldsymbol{x}^{\star}\|$. Hence, in this case we only need to consider all the properties of $h$ in the bounded domain $\{\boldsymbol{x}\in\mathbb{R}^{d}:\|\boldsymbol{x}-\boldsymbol{x}^{\star}\|\leq\|\boldsymbol{x}_{1}-\boldsymbol{x}^{\star}\|\}$. This is well-known via the subgradient projector perspective, as $\boldsymbol{x}_{t}$ is guaranteed to approach the set $\bigcap_{\boldsymbol{\xi}}\{\boldsymbol{x}:h(\boldsymbol{x},\boldsymbol{\xi})=0\}$ at each iteration if it is non-empty. This property was observed in Gower et al. [20] for SPS+ [15], but here it holds more generally. For example, $h(\boldsymbol{x},\boldsymbol{\xi})=(f(\boldsymbol{x},\boldsymbol{\xi})-a)_{+}$, where $a\geq\sup_{\xi}f(\boldsymbol{x}^{\star},\boldsymbol{\xi})$, would also have no neighbourhood of convergence and satisfies the assumption of the theorem.
@@
In the proof of Theorem 3, if one stops before taking expectations, one obtains a regret guarantee on a sequence of arbitrary losses $h(\boldsymbol{x},\boldsymbol{\xi}_{t})$. Such regret scales as the sum of the loss in $\boldsymbol{x}^{\star}$. This is exactly the $L^{\star}$ bound that we mentioned in Section 2. Indeed, this kind of updates and guarantees were already obtained in the online learning literature for the special case of linear predictors by the Passive-Aggressive family of algorithms [11].
Besides covering a number of previous algorithmic variants, we also extend the previous known guarantees. In particular, Loizou et al. [29] only studied SPS  in the non-smooth setting but did not include SPSmax. [^4] Although Loizou et al. [29] state that SPS analysis can be readily extended to SPSmax this does not seem to be the case due to the non-convexity of the min function, as demonstrated by our different proof technique in Appendix C. Theorem 3 shows for the first time that SPSmax  is adaptive to the entire range of upper curvature of the function, from Lipschitz to smooth functions. In Appendix D we also show additional results. Moreover, the second result in the smooth case is new, and it allows to recover the SGD guarantee on $H$ when $\gamma$ is sufficiently small. For example, we include a precise statement for SPS+ when $f(\boldsymbol{x},\boldsymbol{\xi})$ is $L$-self-bounded, that recovers the guarantee in Gower et al. [20, Corollary 2.3].
6 Neighbourhood of Convergence and Instability of the Polyak Stepsize